OpenStudy (anonymous):

What is the solution to the equation 9(x + 1) = 27

OpenStudy (ipwnbunnies):

Give it your best shot, first.

OpenStudy (anonymous):

x = 2.5 x = 0.5 x = -0.5 x = -3.5

OpenStudy (anonymous):

is it A?

OpenStudy (ipwnbunnies):

o-o None of those are correct.

OpenStudy (anonymous):

isnt it supposed to be 2?

OpenStudy (texaschic101):

9(x + 1) = 27 -- distribute the 9 through the parenthesis 9 * x + 9 * 1 = 27 9x + 9 = 27 -- subtract 9 from both sides 9x + 9 - 9 = 27 - 9 9x = 18 x = 2 check. 9(2 + 1) = 27 9 (3) = 27 27 = 27 (correct) yep..its 2

OpenStudy (shamil98):

or instead of doing all of that ... 9(x + 1) = 27 divide both sides by 9.. x+1 = 3 subtract 1 x = 2

OpenStudy (anonymous):

It would actually be 0.5 since the equation is 9^(x + 1) = 27